OpenStudy (anonymous):

please help! 3 − (2x − 5) < −4(x + 2)

OpenStudy (anonymous):

3-(2x-5)<-4(x+2) 3-2x+5<-4x-8 -2x+4x<3-8+5 2x<3-8+5 2x<-5+5 2x<0

OpenStudy (anonymous):

i think thats how you do that

OpenStudy (anonymous):

ok thanks so much

OpenStudy (akashdeepdeb):

3 − (2x − 5) < −4(x + 2) = 3 - 2x + 5 < -4x - 8 = -2x + 2 < -4x - 8 Divide equation by -2, [change the inequality thus] = x - 1 > 2x + 4 = -5 > x Or x < -5 @MexicanHoneyBadger 's solution is not correct.

OpenStudy (anonymous):

cool thanks

OpenStudy (akashdeepdeb):

:)